On April 2, 2017, Montana Mining Co. pays $4,364,740 for an ore deposit containing 1,489,000 tons. The company installs machinery in the mine costing $150,600, with an estimated seven-year life and no salvage value. The machinery will be abandoned when the ore is completely mined. Montana begins mining on May 1, 2017, and mines and sells 184,600 tons of ore during the remaining eight months of 2017.
  
Prepare the December 31, 2017, entries to record both the ore deposit depletion and the mining machinery depreciation. Mining machinery depreciation should be in proportion to the mine’s depletion
Answer #1

Solution

Montana Mining Co

Entries to record both the ore deposit depletion and the mining machinery depreciation:

Date

Account Titles and Explanation

Ref. No.

Debit

Credit

31-Dec-17

Depletion Expense - Mineral Deposit

$541,122

Accumulated Depletion - Mineral Deposit

$541,122

(To record depletion expense)

31-Dec-17

Depreciation Expense - Machinery

$18,670

Accumulated Depreciation - Machinery

$18,670

(To record depreciation on machinery)

Computations:

  1. Depletion expense = rate per ton x tons mined

Rate per ton = cost of mine/estimated tons

Cost of mine = $4,364,740

Estimated tons = 1,489,000

Cost per ton = 4,364,740/1,489,000 = $2.931

Ore mined in 2017 = 184,600 tons

Depletion expense = $2.931 x 184,600 = $541,122

  1. Depreciation expense = machinery rate per ton x tons mined

Machinery cost = $150,600

Machinery rate per ton = $150,600/1,489,000 = $0.101

Ore mined in 2017 = 184,600 tons

Depreciation expense = $0.101 x 184,600 = $18,670

Note: the tons of ore mined is used as basis for depreciation on machinery as the question specifies that mining machinery depreciation should be in the proportion to the mine’s depletion.